OpenStudy (anonymous):

f(x)=5^(x+1) -25 Find x and y intercept

OpenStudy (anonymous):

What is f(0)?

OpenStudy (anonymous):

\[ f(0)= 5^{0+1} -25= 5-25=-20 \]

OpenStudy (anonymous):

For which value of x, f(x) =0?

OpenStudy (anonymous):

\[ 0=5^{x+1}-25\\ 5^{x+1}=25=5^2\\ x+1 =2\\ x=1 \]